Consider a stock that is expected to pay a dividend of $0.93 a year from now. The current price of the stock is $45.18. The expected rate of return on the stock is 13%. What must be the expected growth rate of the dividends? Enter your answer as a percentage. Do not enter the percentage sign into your answer.

Enter your response below rounded to 2 DECIMAL PLACES
Answer #1

Ans 10.94

P0 = Price of Share
D1 = Current Dividend
Ke = Cost of Equity
g = growth rate
P0 = D1 / (Ke - g)
45.18 = 0.93 / (13%- g)
13% - g = 0.93 / 45.18
13% - g = 2.06%
g = 13% - 2.06%
g = 10.94%